OpenStudy (anonymous):

x ^{1/8}*y ^{1/3}*z ^{1/6}=

OpenStudy (mathteacher1729):

What are you supposed to do with this?

OpenStudy (anonymous):

what should we do with this equation?

OpenStudy (anonymous):

you can change all of then in radical ....

OpenStudy (anonymous):

write as a single radical expression I have no Idea where to start

OpenStudy (anonymous):

\[\sqrt[9]{x}*\sqrt[3]{y}*\sqrt[6]{?}\]

OpenStudy (anonymous):

\[\sqrt[6]{x}\sqrt[3]{y}\sqrt[6]{}\]

OpenStudy (anonymous):

does 144sqrtxyz make it one radical?

OpenStudy (anonymous):

\[\sqrt[8]{x}*\sqrt[3]{y}*\sqrt[5]{z} \]

OpenStudy (anonymous):

\[\sqrt[144]{xyz}\]

OpenStudy (anonymous):

yeah it is right the way u did!!!

OpenStudy (anonymous):

Really!!! Oh thank goodness! Thank you so much!!